I need to factor completely this equation:
1000P^3 - 27J^3
It doesn't appear that this can factor

Answer by HyperBrain(694)

Quicknote:
a3-b3=(a-b)(a2+ab+b2)
Observation: 1000P3=(10P)3 and 27J3=(3J)3
So
let a=10P
b=3J
1000P3-27J3=(10P-3J)(100P2+30PJ+9J2)
Power up,
HyperBrain!
Share my blogs>>::<<
here
and here